# Restricted Maximum Likelihood Estimation in Generalized Linear Mixed Models

### v1

@@ 2.1 REML Estimation for Linear Mixed Models
To address the above problem, Patterson and Thompson, (1971, 1974) introduced the idea of restricted maximum likelihood (REML) estimation for LMMs, based on maximizing a restricted log-likelihood function. REML can be derived in a number of ways, and here we review the conditional derivation of Verbyla, (1990) for an LMM of the form
$\boldsymbol{y}=\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u}+\boldsymbol{\varepsilon},\quad\boldsymbol{u}\sim N(\boldsymbol{0},\sigma^{2}\boldsymbol{G}),\quad\boldsymbol{\varepsilon}\sim N(\boldsymbol{0},\sigma^{2}\rho\boldsymbol{R}),$ (4)
@@
$\ell_{\text{LMM},J}\{\boldsymbol{\beta},\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y},\boldsymbol{u}\}$ $=-\frac{1}{2}\left\{\log\det(\sigma^{2}\boldsymbol{G})+\log\det(\sigma^{2}\rho\boldsymbol{R})\right\}-\frac{1}{2\sigma^{2}}\boldsymbol{u}^{\top}\boldsymbol{G}^{-1}\boldsymbol{u}$
$\qquad-\frac{1}{2\sigma^{2}\rho}(\boldsymbol{y}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u})^{\top}\boldsymbol{R}^{-1}(\boldsymbol{y}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u}),$
@@
$\left[\begin{array}[]{c}\boldsymbol{y}_{1}\\[4.30554pt] \boldsymbol{y}_{2}\end{array}\right]\sim N\left(\left[\begin{array}[]{c}\boldsymbol{\beta}\\[6.45831pt] \boldsymbol{0}\end{array}\right],\sigma^{2}\left[\begin{array}[]{cc}\boldsymbol{L}_{1}^{\top}\boldsymbol{V}\boldsymbol{L}_{1}&\boldsymbol{L}_{1}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}\\[4.30554pt] \boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{1}&\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}\end{array}\right]\right),$
@@
where $\boldsymbol{V}=\rho\boldsymbol{R}+\boldsymbol{Z}\boldsymbol{G}\boldsymbol{Z}^{\top}$. It follows that the joint log-likelihood $\ell_{\text{LMM},J}\{\boldsymbol{\beta},\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y},\boldsymbol{u}\}$ can be written as the sum of the two contributions: one from the conditional distribution of $\boldsymbol{y}_{1}$ given $\boldsymbol{y}_{2}$, and one from the marginal distribution of $\boldsymbol{y}_{2}$. That is, $\boldsymbol{y}_{1}\,|\,\boldsymbol{y}_{2}\sim N\big{(}\boldsymbol{\beta}-\boldsymbol{L}_{1}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}(\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2})^{-1}\boldsymbol{y}_{2},\sigma^{2}(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})^{-1}\big{)}$, and $\boldsymbol{y}_{2}\sim N\big{(}\boldsymbol{0},\sigma^{2}\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}\big{)}$. It is the marginal distribution of $\boldsymbol{y}_{2}$ which forms the restricted log-likelihood function
$\ell_{\text{LMM},R}\{\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y}\}$
@@
$=-\frac{1}{2}\left\{(n-p)\log(\sigma^{2})+\log\det(\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2})\right\}-\frac{1}{2\sigma^{2}}\boldsymbol{y}_{2}^{\top}(\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2})^{-1}\boldsymbol{y}_{2}$
$=-\frac{1}{2}\left\{(n-p)\log(\sigma^{2})+\log\det(\boldsymbol{V})\right\}-\frac{1}{2\sigma^{2}}\boldsymbol{y}^{\top}\boldsymbol{P}\boldsymbol{y}-\frac{1}{2}\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X}),$ (5)
@@
where $\boldsymbol{P}=\boldsymbol{V}^{-1}-\boldsymbol{V}^{-1}\boldsymbol{X}\big{(}\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X}\big{)}^{-1}\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}$ and constants are again omitted. Note expression (5) does not depend on the fixed effect coefficients $\boldsymbol{\beta}$. The final term $-(1/2)\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})$ is sometimes known in the literature as the “REML adjustment”, and forms similar to this will appear throughout this review. Maximization of $\ell_{\text{LMM},R}\{\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y}\}$ leads to REML estimators of the random effects covariance matrix and dispersion parameter. It can be shown that REML estimators have the properties of likelihood-based estimators, in that they are asymptotically normally distributed with covariance matrix based on the information matrix derived from $\ell_{\text{LMM},R}\{\phi,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y}\}$; see Verbyla, (1990) and Cressie and Lahiri, (1993) for details.
@@
The conditional derivation above represents one of several alternative derivations by which REML estimators can be derived for LMMs. Importantly, all formulations of REML for LMMs produce the same objective function as that given in equation (5). By contrast, this variety of derivations has inspired a multitude of REML likelihood functions and estimators in GLMMs. We review these in the next four sections.
@@ 3 Approximate Linearization
The greater part of the research on REML estimation for GLMMs uses an approximate linearization strategy. Such procedures involve replacing the GLMM by an approximate linear mixed model, after which various LMM results for constructing REML estimates can be adapted. We choose to review approximate linearization methods through three sets of representative, connected papers: Schall, (1991) and Wolfinger and O’Connell, (1993) who obtained REML estimates by applying results on an adjusted dependent variable, McGilchrist, (1994) whose computation of REML estimates is framed within an iterative scheme for calculating best linear unbiased predictors, and Breslow and Clayton, (1993) and Engel and Keen, (1994) who proposed quasi-likelihood estimation methods for GLMMs.
@@ 3.1 Linearization and Adjusted Dependent Variables
Schall, (1991) considered a GLMM of the form $g(\boldsymbol{\mu})=\boldsymbol{X}\boldsymbol{\beta}+\sum_{k=1}^{K}\boldsymbol{Z}_{k}\boldsymbol{u}_{k}$ and $\mbox{Cov}(\boldsymbol{u})=\boldsymbol{D}(\boldsymbol{\theta})=\mbox{diag}(\sigma^{2}_{1}\boldsymbol{I}_{q_{1}},\ldots,\sigma^{2}_{K}\boldsymbol{I}_{q_{K}})$, where $\boldsymbol{\theta}=(\sigma^{2}_{1},\ldots,\sigma^{2}_{K})^{\top}$ are the variance components. This model is a special case of (1) where the random effects model matrix comprises a set of $k=1,\ldots,K$ sparse submatrices $\boldsymbol{Z}_{k}$ of dimension $n\times q_{k}$, with $q=\sum_{k=1}^{K}q_{k}$, while the corresponding random effects $\boldsymbol{u}=(\boldsymbol{u}_{1}^{\top},\ldots,\boldsymbol{u}^{\top}_{K})^{\top}$ are structured such that the $q_{k}$ elements of $\boldsymbol{u}_{k}$ are drawn independently from a random effects distribution with mean zero and variance component $\sigma^{2}_{k}$. This particular form of GLMM is often associated with nested/multi-level data where only random intercepts are included, as well as data from crossed-effect designs. We remark that the original formulation in Schall, (1991) makes assumptions only on the expectation and variance of the random effects distribution, although the results shown below assume normality for $\boldsymbol{u}$.
The method of Schall, (1991) uses an approximate linearization of the link function $g(y_{i})\approx\eta_{i}+g^{\prime}(\mu_{i})(y_{i}-\mu_{i})\triangleq\xi_{i}$, where $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})^{\top}$ denotes a vector of adjusted dependent variables (also known as working responses; McCullagh and Nelder,, 1989). Note this linearization is commonly used in the context of estimation generalized linear and generalized additive models (via iterative reweighted least squares; Wood,, 2017). Based on this adjusted dependent variable we can construct an approximate linearized GLMM as
@@
where $\boldsymbol{W}^{-1}$ is a diagonal weight matrix with diagonal elements $\text{Var}\{g^{\prime}(\mu_{i})(y_{i}-\mu_{i})|\boldsymbol{u}\}=\phi V(\mu_{i})\{g^{\prime}(\mu_{i})\}^{2}$. Regarding equation (6) as an LMM with marginal mean and covariance $\mbox{E}(\boldsymbol{\xi}|\boldsymbol{u})=\boldsymbol{X}\boldsymbol{\beta}$ and $\mbox{Cov}(\boldsymbol{\xi})=\boldsymbol{W}^{-1}+\boldsymbol{Z}\boldsymbol{D}(\boldsymbol{\theta})\boldsymbol{Z}^{\top}$, we can then apply established methods for computing REML estimates in LMMs. In particular, Schall, (1991) proposed a two-step iterative procedure adapted from Fellner, (1986, 1987), which itself is built upon the work of Henderson, (1963) and Harville, (1977). First, given current values of all the parameters $(\boldsymbol{\beta}^{\top},\boldsymbol{\theta}^{\top},\phi)^{\top}$ and random effects $\boldsymbol{u}$, update the adjusted dependent variable $\boldsymbol{\xi}^{\dagger}=\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u}$, and compute $\boldsymbol{D}(\boldsymbol{\theta})$ and $\boldsymbol{W}$. Then, given these updated values, compute $\boldsymbol{\beta}$ and $\boldsymbol{u}$ by solving the mixed model equations of Henderson, (1963),
@@
$\boldsymbol{A}\left[\begin{array}[]{c}\boldsymbol{\beta}\\[4.30554pt] \boldsymbol{u}\end{array}\right]=\left[\begin{array}[]{cc}\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{X}&\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{Z}\\[4.30554pt] \boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{X}&\boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{Z}+\boldsymbol{D}(\boldsymbol{\theta})^{-1}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{\beta}\\[4.30554pt] \boldsymbol{u}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{\xi}^{\dagger}\\[4.30554pt] \boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{\xi}^{\dagger}\end{array}\right].$ (7)
@@
Finally, to update the REML estimates of the variance components, let $\boldsymbol{T}=\{\boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{Z}-\boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{Z}+\boldsymbol{D}(\boldsymbol{\theta})^{-1}\}^{-1}$ denote the matrix formed by the last $q$ rows and $q$ columns of $\boldsymbol{A}^{-1}$. Partition $\boldsymbol{T}$ into blocks $\boldsymbol{T}_{kk^{\prime}}$ for $k,k^{\prime}=1,\ldots,K$ conformably with the diagonal structure of $\boldsymbol{D}(\boldsymbol{\theta})$, and update the elements of $\boldsymbol{\theta}$ via the fixed-point formula $\sigma^{2}_{k}\Leftarrow(\boldsymbol{u}_{k}^{\top}\boldsymbol{u}_{k})/\{q_{k}-\sigma^{-2}_{k}\mbox{trace}(\boldsymbol{T}_{kk})\}$. If the dispersion parameter $\phi$ also requires estimation, a further step can be taken again using a REML-like estimator; see Fellner, (1986) for more details.
@@
Schall, (1991) pointed out that the proposed approach to REML estimation produces the same estimates of the variance components as the algorithm proposed by Stiratelli et al., (1984) using the integrated likelihood approach, which we review in Section 4. In the special case of a binomial GLMM using a probit link and single set of random effects, i.e., $K=1$, the approximate linearization approach of Schall, (1991) also corresponds to the joint maximization method of Gilmour et al., (1985).
@@
Wolfinger, (1993) extended the work of Schall, (1991) to construct REML estimates for nonlinear mixed models, but of more relevance here is the work of Wolfinger and O’Connell, (1993), who studied GLMMs in the form given by (1), i.e., allowing for more general structures in $\boldsymbol{D}(\boldsymbol{\theta})$ compared to Schall, (1991). Specifically, Wolfinger and O’Connell, (1993) arrived at the approximate linearized GLMM (6), but then adapted the work of Lindstrom and Bates, (1990) and explicitly assumed a multivariate normal distribution for $\boldsymbol{e}$. That is, their strategy is to directly specify $\boldsymbol{\xi}\sim N(\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u},\boldsymbol{W}^{-1})$ and then, analogously to Section 2.1 except that $\phi$ is also profiled, obtain at the restricted log-likelihood function
$\ell_{\text{lin},R}(\boldsymbol{\theta};\boldsymbol{y})=-\frac{1}{2}\log\det(\boldsymbol{V})-\frac{n-p}{2}\log\boldsymbol{r}^{\top}\boldsymbol{V}^{-1}\boldsymbol{r}-\frac{1}{2}\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X}),$ (8)
@@
where $\boldsymbol{V}=\boldsymbol{W}^{-1}+\boldsymbol{Z}\boldsymbol{D}(\boldsymbol{\theta})\boldsymbol{Z}^{\top}$, $\boldsymbol{r}=\boldsymbol{\xi}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{\xi}$, and constants with respect to the model parameters are omitted. Note the $\boldsymbol{V}$ matrix here is also a function of $\boldsymbol{\beta}$ and $\boldsymbol{u}$ through $\boldsymbol{W}$, unlike in LMMs. Wolfinger and O’Connell, (1993) then iterated between solving the mixed model equations in (7) to update $\boldsymbol{\beta}$ and $\boldsymbol{u}$ plus using a profile likelihood estimate of $\phi$, and maximizing $\ell_{\text{lin},R}(\boldsymbol{\theta})$ to update $\boldsymbol{\theta}$. More generally, equation (8) may be regarded as a GLMM analogue of the restricted log-likelihood function in (5), achieved through approximate linearization, where the term $-(1/2)\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})$ is a REML adjustment.

### v2

@@ 2.1 REML Estimation for Linear Mixed Models
To address the above problem, Patterson and Thompson (1971, 1974) introduced the idea of REML estimation for LMMs, based on maximizing a restricted log-likelihood function. REML can be derived in a number of ways, and here we review the conditional derivation of Verbyla (1990) for an LMM of the form
$\begin{array}[]{c}\boldsymbol{y}=\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u}+\boldsymbol{\varepsilon},\quad\boldsymbol{u}\sim N(\boldsymbol{0},\sigma^{2}\boldsymbol{G}),\\[6.45831pt] \boldsymbol{\varepsilon}\sim N(\boldsymbol{0},\sigma^{2}\rho\boldsymbol{R}),\end{array}$ (4)
@@
$\ell_{\text{LMM},J}\{\boldsymbol{\beta},\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y},\boldsymbol{u}\}=-\frac{1}{2}\left\{\log\det(\sigma^{2}\boldsymbol{G})\right\}$
$-\frac{1}{2}\left\{\log\det(\sigma^{2}\rho\boldsymbol{R})\right\}-\frac{1}{2\sigma^{2}}\boldsymbol{u}^{\top}\boldsymbol{G}^{-1}\boldsymbol{u}$
$-\frac{1}{2\sigma^{2}\rho}(\boldsymbol{y}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u})^{\top}\boldsymbol{R}^{-1}(\boldsymbol{y}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u}),$
@@
$\left[\begin{array}[]{c}\boldsymbol{y}_{1}\\[4.30554pt] \boldsymbol{y}_{2}\end{array}\right]\sim N\left(\left[\begin{array}[]{c}\boldsymbol{\beta}\\[6.45831pt] \boldsymbol{0}_{n-p}\end{array}\right],\sigma^{2}\left[\begin{array}[]{cc}\boldsymbol{L}_{1}^{\top}\boldsymbol{V}\boldsymbol{L}_{1}&\boldsymbol{L}_{1}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}\\[4.30554pt] \boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{1}&\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}\end{array}\right]\right),$
@@
where $\boldsymbol{V}=\rho\boldsymbol{R}+\boldsymbol{Z}\boldsymbol{G}\boldsymbol{Z}^{\top}$ and $\boldsymbol{0}_{n-p}$ is an $n-p$ vector of zeros. It follows that the joint log-likelihood $\ell_{\text{LMM},J}\{\boldsymbol{\beta},\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y},\boldsymbol{u}\}$ can be written as the sum of the two contributions: one from the conditional distribution of $\boldsymbol{y}_{1}$ given $\boldsymbol{y}_{2}$, and one from the marginal distribution of $\boldsymbol{y}_{2}$. That is,
$\begin{array}[]{c}\boldsymbol{y}_{1}\,|\,\boldsymbol{y}_{2}\sim N\big{(}\boldsymbol{\beta}-\boldsymbol{L}_{1}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}(\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2})^{-1}\boldsymbol{y}_{2},\\[4.30554pt] \qquad\qquad\sigma^{2}(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})^{-1}\big{)}\\[6.45831pt] \mbox{and}\quad\boldsymbol{y}_{2}\sim N\big{(}\boldsymbol{0},\sigma^{2}\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2}\big{)}.\end{array}$
It is the marginal distribution of $\boldsymbol{y}_{2}$ which forms the restricted log-likelihood function
@@
$\begin{split}&\ell_{\text{LMM},R}\{\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y}\}=-\frac{1}{2}\Big{\{}(n-p)\log(\sigma^{2})\\ &\quad+\log\det(\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2})\Big{\}}-\frac{1}{2\sigma^{2}}\boldsymbol{y}_{2}^{\top}(\boldsymbol{L}_{2}^{\top}\boldsymbol{V}\boldsymbol{L}_{2})^{-1}\boldsymbol{y}_{2}\\ &=-\frac{1}{2}\left\{(n-p)\log(\sigma^{2})+\log\det(\boldsymbol{V})\right\}\\ &\quad-\frac{1}{2\sigma^{2}}\boldsymbol{y}^{\top}\boldsymbol{P}\boldsymbol{y}-\frac{1}{2}\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X}),\end{split}$ (5)
@@
where $\boldsymbol{P}=\boldsymbol{V}^{-1}-\boldsymbol{V}^{-1}\boldsymbol{X}\big{(}\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X}\big{)}^{-1}\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}$ and constants are again omitted. Note expression (5) does not depend on the fixed effect coefficients $\boldsymbol{\beta}$. The final term $-(1/2)\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})$ is sometimes known in the literature as the “REML adjustment”, and forms similar to this will appear throughout this review. Maximization of $\ell_{\text{LMM},R}\{\rho,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y}\}$ leads to REML estimators of the random effects covariance matrix and dispersion parameter. It can be shown that REML estimators have the properties of likelihood-based estimators, in that they are asymptotically normally distributed with covariance matrix based on the information matrix derived from $\ell_{\text{LMM},R}\{\phi,\boldsymbol{\theta},\text{vech}(\boldsymbol{R});\boldsymbol{y}\}$, and assuming here the dimension of $\boldsymbol{\beta}$ ($p$) does not grow with sample size ($n$); see Verbyla (1990) and Cressie and Lahiri (1993) for details.
@@
The conditional derivation above represents one of several alternative derivations by which REML estimators can be derived for LMMs. Importantly, all formulations of REML for LMMs produce the same objective function as that given in equation (5). By contrast, this variety of derivations has inspired a multitude of REML likelihood functions and estimators in GLMMs. We review these in the next five sections.
@@ 3 Approximate Linearization
The greater part of the research on REML estimation for GLMMs uses an approximate linearization strategy. Such procedures involve replacing the GLMM by an approximate linear mixed model, after which various LMM results for constructing REML estimates can be adapted. We choose to review approximate linearization methods through three sets of representative, connected papers: Schall (1991) and Wolfinger and O’Connell (1993) who obtained REML estimates by applying results on an adjusted dependent variable, McGilchrist (1994) whose computation of REML estimates is framed within an iterative scheme for calculating best linear unbiased predictors, and Breslow and Clayton (1993) and Engel and Keen (1994) who proposed quasi-likelihood estimation methods for GLMMs.
@@ 3.1 Linearization and Adjusted Dependent Variables
Schall (1991) considered a GLMM of the form
$\begin{array}[]{c}g(\boldsymbol{\mu})=\boldsymbol{X}\boldsymbol{\beta}+\sum_{k=1}^{K}\boldsymbol{Z}_{k}\boldsymbol{u}_{k},\\[6.45831pt] \mbox{Cov}(\boldsymbol{u})=\boldsymbol{D}(\boldsymbol{\theta})=\mbox{diag}(\sigma^{2}_{1}\boldsymbol{I}_{q_{1}},\ldots,\sigma^{2}_{K}\boldsymbol{I}_{q_{K}}),\end{array}$
where $\boldsymbol{\theta}=(\sigma^{2}_{1},\ldots,\sigma^{2}_{K})^{\top}$ are the variance components. This model is a special case of (1) where the random effects model matrix comprises a set of $k=1,\ldots,K$ sparse submatrices $\boldsymbol{Z}_{k}$ of dimension $n\times q_{k}$, with $q=\sum_{k=1}^{K}q_{k}$, while the corresponding random effects $\boldsymbol{u}=(\boldsymbol{u}_{1}^{\top},\ldots,\boldsymbol{u}^{\top}_{K})^{\top}$ are structured such that the $q_{k}$ elements of $\boldsymbol{u}_{k}$ are drawn independently from a random effects distribution with mean zero and variance component $\sigma^{2}_{k}$. This particular form of GLMM is often associated with nested/multi-level data where only random intercepts are included, as well as data from crossed-effect designs. We remark that the original formulation in Schall (1991) makes assumptions only on the expectation and variance of the random effects distribution, although the results shown below assume normality for $\boldsymbol{u}$. More generally, the approach of Schall (1991) which we describe below only applies to GLMMs with independent random effects, and strictly to cases where the random effects precision matrix is linear in the variance parameters (see also Section 6.4 of Wood, 2017).
The method of Schall (1991) uses an approximate linearization of the link function $g(y_{i})\approx\eta_{i}+g^{\prime}(\mu_{i})(y_{i}-\mu_{i})\triangleq\xi_{i}$, where $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})^{\top}$ denotes a vector of adjusted dependent variables (also known as working responses; McCullagh and Nelder, 1989). Note this linearization is commonly used in the context of estimation for generalized linear and generalized additive models (via iterative reweighted least squares; Wood, 2017). Based on this adjusted dependent variable we can construct an approximate linearized GLMM as
@@
where $\boldsymbol{W}^{-1}$ is a diagonal weight matrix with diagonal elements $\text{Var}\{g^{\prime}(\mu_{i})(y_{i}-\mu_{i})|\boldsymbol{u}\}=\phi V(\mu_{i})\{g^{\prime}(\mu_{i})\}^{2}$. Regarding the expressions in (6) as an LMM with marginal mean and covariance $\mbox{E}(\boldsymbol{\xi}|\boldsymbol{u})=\boldsymbol{X}\boldsymbol{\beta}$ and $\mbox{Cov}(\boldsymbol{\xi})=\boldsymbol{W}^{-1}+\boldsymbol{Z}\boldsymbol{D}(\boldsymbol{\theta})\boldsymbol{Z}^{\top}$, we can then apply established methods for computing REML estimates in LMMs. In particular, Schall (1991) proposed a two-step iterative procedure adapted from Fellner (1986, 1987), which itself is built upon the work of Henderson (1963) and Harville (1977). First, given current values of all the parameters $(\boldsymbol{\beta}^{\top},\phi,\boldsymbol{\theta}^{\top})^{\top}$ and random effects $\boldsymbol{u}$, update the adjusted dependent variable $\boldsymbol{\xi}^{\dagger}=\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u}$, and compute $\boldsymbol{D}(\boldsymbol{\theta})$ and $\boldsymbol{W}$. Then, given these updated values, compute $\boldsymbol{\beta}$ and $\boldsymbol{u}$ by solving the mixed model equations of Henderson (1963),
@@
$\begin{split}\boldsymbol{A}\left[\begin{array}[]{c}\boldsymbol{\beta}\\[4.30554pt] \boldsymbol{u}\end{array}\right]&=\left[\begin{array}[]{cc}\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{X}&\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{Z}\\[4.30554pt] \boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{X}&\boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{Z}+\boldsymbol{D}(\boldsymbol{\theta})^{-1}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{\beta}\\[4.30554pt] \boldsymbol{u}\end{array}\right]\\ &=\left[\begin{array}[]{c}\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{\xi}^{\dagger}\\[4.30554pt] \boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{\xi}^{\dagger}\end{array}\right].\end{split}$ (7)
Finally, to update the REML estimates of the variance components, let
$\boldsymbol{T}$ =$=\{\boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{Z}$
$\,\,\quad-\boldsymbol{Z}^{\top}\boldsymbol{W}\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{W}\boldsymbol{Z}+\boldsymbol{D}(\boldsymbol{\theta})^{-1}\}^{-1}$
@@
denote the matrix formed by the last $q$ rows and $q$ columns of $\boldsymbol{A}^{-1}$. Partition $\boldsymbol{T}$ into blocks $\boldsymbol{T}_{kk^{\prime}}$ for $k,k^{\prime}=1,\ldots,K$ conformably with the diagonal structure of $\boldsymbol{D}(\boldsymbol{\theta})$, and update the elements of $\boldsymbol{\theta}$ via the fixed-point formula $\sigma^{2}_{k}\Leftarrow(\boldsymbol{u}_{k}^{\top}\boldsymbol{u}_{k})/\{q_{k}-\sigma^{-2}_{k}\mbox{trace}(\boldsymbol{T}_{kk})\}$. If the dispersion parameter $\phi$ also requires estimation, a further step can be taken again using a REML-like estimator; see Fellner (1986) for more details.
@@
Schall (1991) pointed out that the proposed approach to REML estimation produces the same estimates of the variance components as the algorithm proposed by Stiratelli, Laird and Ware (1984) using the integrated likelihood approach, which we review in Section 4. It is also worth noticing that the method of Schall (1991) can be obtained simply from the derivative of the Laplace-approximate marginal log-likelihood computed with respect to the fixed and random effects; see Section 6.4.1 of Wood (2017) for further details as well as additional connections to the penalized quasi-likelihood approach reviewed in Section 3.3. In the special case of a binomial GLMM using a probit link and single set of random effects, i.e., $K=1$, the approximate linearization approach of Schall (1991) also corresponds to the joint maximization method of Gilmour, Anderson and Rae (1985).
@@
Wolfinger (1993) extended the work of Schall (1991) to construct REML estimates for nonlinear mixed models, although Wolfinger’s approach applies the Laplace approximation to the marginal log-likelihood function of the nonlinear mixed model along with an improved first-order approximation to the nonlinear function of Lindstrom and Bates (1990); a REML adjustment akin to the final log-determinant term of (5) for estimating the variance components is then incorporated. Of more relevance here is the work of Wolfinger and O’Connell (1993), who studied GLMMs in the form given by (1), i.e., allowing for more general structures in $\boldsymbol{D}(\boldsymbol{\theta})$ compared to Schall (1991). Specifically, Wolfinger and O’Connell (1993) arrived at the approximate linearized GLMM (6), but then adapted the work of Lindstrom and Bates (1990) and explicitly assumed a multivariate normal distribution for $\boldsymbol{e}$. That is, their strategy is to directly specify $\boldsymbol{\xi}\sim N(\boldsymbol{X}\boldsymbol{\beta}+\boldsymbol{Z}\boldsymbol{u},\boldsymbol{W}^{-1})$ and then, analogously to Section 2.1 except that $\phi$ is also profiled, obtain the restricted log-likelihood function
$\begin{split}\ell_{\text{lin},R}(\boldsymbol{\theta};\boldsymbol{y})&=-\frac{1}{2}\log\det(\boldsymbol{V})-\frac{n-p}{2}\log\boldsymbol{r}^{\top}\boldsymbol{V}^{-1}\boldsymbol{r}\\ &\quad-\frac{1}{2}\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X}),\end{split}$ (8)
@@
where $\boldsymbol{r}=\boldsymbol{\xi}-\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{\xi}$, $\boldsymbol{V}=\boldsymbol{W}^{-1}+\boldsymbol{Z}\boldsymbol{D}(\boldsymbol{\theta})\boldsymbol{Z}^{\top}$, and constants with respect to the model parameters are omitted. Note the $\boldsymbol{V}$ matrix here is also a function of $\boldsymbol{\beta}$ and $\boldsymbol{u}$ through $\boldsymbol{W}$, unlike in LMMs. Wolfinger and O’Connell (1993) then iterated between solving the mixed model equations in (7) to update $\boldsymbol{\beta}$ and $\boldsymbol{u}$ plus using a profile likelihood estimate of $\phi$, and maximizing $\ell_{\text{lin},R}(\boldsymbol{\theta})$ to update $\boldsymbol{\theta}$. More generally, equation (8) may be regarded as a GLMM analogue of the restricted log-likelihood function in (5), achieved through approximate linearization, where the term $-(1/2)\log\det(\boldsymbol{X}^{\top}\boldsymbol{V}^{-1}\boldsymbol{X})$ is a REML adjustment.